# Banach algebras associated to twisted étale groupoids: simplicity and pure infiniteness

### v2

@@ 7.3. Tight algebras of inverse semigroups
In this subsection we define reduced and essential tight $L^{P}$-operator algebras associated to inverse semigroups, and phrase our main results in terms of the inverse semigroup. This extends, and sometimes improves, the corresponding results for $C^{*}$-algebras.
We fix an inverse semigroup $S$ with zero 00 ($S$ is not neccessarily unital). The set of idempotents $E:=E(S)$ is a semilattice with the ‘meet’ and ‘preorder’ defined by $e\wedge f:=ef$ and $e\leq f$ if and only if $ef=e$. The set of filters $\widehat{E}$ in $E$, with the product topology inherited from the Cantor space $\{0,1\}^{E}$, is a totally disconnected Hausdorff space called the spectrum of $E$. A cover of $e\in E$ is a finite set $F\subseteq E$ such that for every non-zero $z\leq e$ we have $zf\neq 0$ for some $f\in F$. A filter $\phi\in\hat{E}$ is tight if for every $e\in\phi$ and any cover $F\subseteq eE$ of $e$ we have $\phi\cap F\neq\emptyset$. The set of tight filters $\widehat{E}_{\operatorname{tight}}$ is called the tight spectrum of $E$. It is the closure in $\hat{E}$ of the set of all ultrafilters (if $E$ is a Boolean algebra then the tight filters are precisely the ultrafilters). For each $e\in E$ the set $Z_{e}:=\{\phi\in\widehat{E}_{\operatorname{tight}}:e\in\phi\}$ is compact open in $\widehat{E}_{\operatorname{tight}}$. For each $t\in S$ we have a homeomorphism $h_{t}:Z_{t^{*}t}\to Z_{tt^{*}}$ given by
@@
Let $u=\{u(s,t)\}_{s,t\in S}$ be a twist of the action $h:S\to\operatorname{PHomeo}(\widehat{E}_{{\operatorname{tight}}})$ described above. That is, $u(s,t)\in C_{u}(Z_{t^{*}s^{*}st})$, $t,s\in S$, are maps satisfying conditions in [BE11, Definition 4.1], or in [BKM23, Definition 3.6], and let $\mathcal{L}_{u}$ be the twist associated to $u$, as in [BKM23, Subsection 4.3]. Then $(\mathcal{G}_{{\operatorname{tight}}}(S),\mathcal{L}_{u})$ is an ample twisted groupoid, and every ample twisted groupoid is of this form, see Remark 7.25 below.
We will now represent the inverse semigroup $S$, taking into account the twist $u$, via spatial partial isometries on the spaces $\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))$, $p\in[1,\infty]$. We recall that for any localizable measure $\mu$ (and so in particular for a counting measure) we have a natural inverse semigroup $\operatorname{SPIso}(L^{p}(\mu))\subseteq B(L^{p}(\mu))$ of contractive weighted composition operators called spatial partial isometries on $L^{p}(\mu)$, see [BKM23, Definition 2.25], [Phi12, Definition 6.4], [Gar21, Definition 6.2].
@@
Let $S$ be an inverse semigroup with zero and let $u=\{u(s,t)\}_{s,t\in S}$ be a twist of the associated action $h:S\to\operatorname{PHomeo}(\widehat{E}_{{\operatorname{tight}}})$. For each $p\in[1,\infty]$ we have a map $v:S\to\operatorname{SPIso}\big{(}\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))\big{)}$ into spatial isometries on $\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))$ given by
@@
where $\xi\in\ell^{p}(\mathcal{G})$, $\phi\in Z_{s^{*}s}$, $t,s\in S$. Denoting by $v_{t}^{*}$ the unique generalized inverse of $v_{t}$ in $\operatorname{SPIso}\big{(}\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))\big{)}$ we get that the space
@@
is in fact a Banach algebra naturally isometrically isomorphic to the twisted reduced groupoid algebra $F_{\operatorname{r}}^{p}(\mathcal{G}_{{\operatorname{tight}}}(S),\mathcal{L}_{u})$ via the isomorphism that restricts to $\operatorname{\overline{span}}\{v_{t}v_{t}^{*}:t\in S\}\cong C_{0}(\widehat{E}_{\operatorname{tight}})$.
@@
By [BKM23, Remark 5.2] we have an isometric representation $\pi:C_{0}(\widehat{E}_{\operatorname{tight}})\to B\big{(}\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))\big{)}$ given by the formula $\pi(a)\xi[s,x]=a\big{(}h_{s}(x)\big{)}\xi([s,x])$ and then the pair $(\pi,v)$ forms a covariant representation of the twisted action that integrates to the regular representation of $F(\mathcal{G}_{{\operatorname{tight}}}(S),\mathcal{L}_{u})$ on $\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))$. In particular, $v_{t}$ is a spatial partial isometry, with the generalized inverse being $v_{t}^{*}=v_{t^{*}}\pi(\overline{u(t,t^{*})})$, and $v_{t}v_{t}^{*}=\pi(1_{Z_{tt^{*}}})$ for each $t\in S$. This implies that we have an isometric isomorphism $C_{0}(\widehat{E}_{\operatorname{tight}})\cong\pi(C_{0}(\widehat{E}_{\operatorname{tight}}))=\operatorname{\overline{span}}\{v_{t}v_{t}^{*}:t\in S\}$ that extends to an isometric isomorphism $F_{\operatorname{r}}^{p}(\mathcal{G}_{{\operatorname{tight}}}(S),\mathcal{L}_{u})\cong F^{p}_{\operatorname{r},{\operatorname{tight}}}(S,u)$ induced by the integrated form $\pi\rtimes v$ of the covariant representation $(\pi,v)$. ∎
@@
When the twist $u$ is trivial we write $F^{p}_{\operatorname{r},{\operatorname{tight}}}(S)$ for $F^{p}_{\operatorname{r},{\operatorname{tight}}}(S,u)$. Then the map $v$ in Proposition 7.23 is a tight representation of $S$ in the sense of [BKM23, Definition 6.12]. In particular, $F^{p}_{\operatorname{r},{\operatorname{tight}}}(S)$ is a reduced version of the tight $L^{p}$-operator algebra $F^{p}_{\operatorname{tight}}(S)$ of $S$ introduced in [BKM23, Definition 6.14]: $v$ induces a natural representation $F^{p}_{\operatorname{tight}}(S)\to F^{p}_{\operatorname{r},{\operatorname{tight}}}(S)$ (which, by [GL17, Theorem 6.19], is an isometric isomorphism if $S$ is countable and the groupoid $\mathcal{G}_{{\operatorname{tight}}}(S)$ is amenable).
@@
By [BKM23, Corollary 6.18 and Lemma 4.23] every twisted ample groupoid $(\mathcal{G},\mathcal{L})$ is isomorphic to $(\mathcal{G}_{{\operatorname{tight}}}(S),\mathcal{L}_{u})$ for some inverse semigroup $S$ (one may take $S$ to be the collection of all compact open bisection in $S(\mathcal{L})$) and some twist $u$. Thus the algebras $F^{p}_{\operatorname{r},{\operatorname{tight}}}(S,u)$ introduced in Proposition 7.23 model all reduced twisted groupoid $L^{p}$-operator algebras associated to ample groupoids.
@@
Retain the notation and assumptions from Proposition 7.23 and fix $p\in[1,\infty]$. Denote by $\mathcal{G}_{\operatorname{H}}(S)$ the set of Hausdorff points in $\mathcal{G}_{{\operatorname{tight}}}(S)$. Then the obvious contractive projection $P_{\operatorname{H}}$ from $\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))$ onto $\ell^{p}(\mathcal{G}_{\operatorname{H}}(S))\subseteq\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))$ commutes with elements of $F^{p}_{\operatorname{r},{\operatorname{tight}}}(S,u)$, and if $S$ is countable then the compressed Banach algebra
@@
Let $P\subseteq[1,\infty]$ be non-empty set. We denote by $F^{P}_{\operatorname{r},{\operatorname{tight}}}(S,u)$ the Banach algebra generated by the $\ell^{\infty}$-direct sum of representations $S\to\operatorname{SPIso}\big{(}\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))\big{)}$, $p\in P$, from Proposition 7.23. We denote by $F^{P}_{\operatorname{e},{\operatorname{tight}}}(S,u)$ the compression of $F^{P}_{\operatorname{r},{\operatorname{tight}}}(S,u)$ to the subspace $\bigoplus_{p\in P}\ell^{p}(\mathcal{G}_{\operatorname{H}}(S))$. We also write $F^{P}_{{\operatorname{tight}}}(S,u):=F^{P}(\mathcal{G}_{{\operatorname{tight}}}(S),\mathcal{L}_{u})$ and skip writing $u$ when $u$ is trivial.
@@
In Definition 7.28 we slightly changed the terminology from [EP16]. Namely, the authors of [EP16] call weakly-fixed what we call fixed, and they call fixed what we call trivially fixed (our naming is consistent with [EP16, Definition 4.1] though). Local contractivness is as in [EP16, Definition 6.1], closedness and minimality apear as conditions in [EP16, Theorems 3.16 and 5.5], respectively. Topological freeness is our invention and it improves (is weaker than) a condition appearing in [EP16, Theorem 4.10], see Remark 7.32 below.
@@
Let $S$ be an inverse semigroup with zero. Then $S$ is closed (resp. topologically free or minimal) if and only if $\mathcal{G}_{\operatorname{tight}}(S)$ is Hausdorff (resp. topologically free or minimal). If $S$ is locally contractive then $\mathcal{G}_{\operatorname{tight}}(S)$ is.
@@
The claims about $S$ being closed, minimal, and locally contractive follow by [EP16, Theorems 3.16, 5.5 and 6.5]. For the claim about topological freeness, recall that topological freeness of $\mathcal{G}_{\operatorname{tight}}(S)$ is equivalent to topological freeness of $h$ in the sense of [KM21, Definition 2.20], by [KM21, Lemma 2.23] This means that for every $t\in S$, the set $\{\phi\in Z_{t^{*}t}:h_{t}(\phi)=\phi\}\setminus\bigcup_{f\in F_{t}}Z_{f}$ has empty interior. By [EP16, Proposition 2.5] and density of ultrafilters in $\widehat{E}_{\operatorname{tight}}$ every non-empty open set in $\widehat{E}_{\operatorname{tight}}$ contains a set of the form $Z_{e}$, for some $e\in E$. Thus topological freeness of $\mathcal{G}_{\operatorname{tight}}(S)$ is equivalent to the condition that for any $t\in S$ and $e\in E$, if $Z_{e}\subseteq\{\phi\in Z_{t^{*}t}:h_{t}(\phi)=\phi\}$, then $Z_{e}\cap\bigcup_{f\in F_{t}}Z_{f}\neq\emptyset$. In view of Lemma 7.30 and the fact that $Z_{e}\cap Z_{f}\neq\emptyset$ if and only if $ef\neq 0$, this in turn is equivalent to saying that for any $t\in S$ and $e\in E$, if $e$ is fixed by $t$, then $ef\neq 0$ for some $f\in F_{t}$, which is exactly what we called topological freeness of $S$. ∎
@@
The condition (ii) in [EP16, Theorem 4.10] tries to characterize effectiveness of $\mathcal{G}_{\operatorname{tight}}(S)$ in terms of $S$. It says that for every $e\in E$ fixed by $t\in S$, $e$ can be covered by elements in $F_{t}$. When $\mathcal{G}_{\operatorname{tight}}(S)$ is Hausdorff (equivalently $S$ is closed), this condition is equivalent to $S$ being topologically free in our sense. This agrees with the fact that effectiveness is equivalent to topological freeness for Hausdorff groupoids (for non-Hausdorff it is in general stronger).

### v3

@@ 7.3. Tight algebras of inverse semigroups
In this subsection, we define reduced and essential tight $L^{P}$-operator algebras associated to inverse semigroups, and phrase our main results in terms of the inverse semigroup. This extends, and sometimes improves, the corresponding results for $C^{*}$-algebras.
We fix an inverse semigroup $S$ with zero 0. We recall our notation from [BKM25, Section 6] for the tight groupoid associated to $S$ in [Exe08]. The set of idempotents $E:=E(S)$ is a semilattice with the ‘meet’ and ‘preorder’ defined by $e\wedge f:=ef$ and $e\leq f$ if and only if $ef=e$. The set of filters $\widehat{E}$ in $E$, with the product topology inherited from the Cantor space $\{0,1\}^{E}$, is a totally disconnected Hausdorff space. A cover of $e\in E$ is a finite set $F\subseteq E$ such that, for every non-zero $z\leq e$, we have $zf\neq 0$ for some $f\in F$. A filter $\phi\in\hat{E}$ is tight if for every $e\in\phi$ and any cover $F\subseteq eE$ of $e$, we have $\phi\cap F\neq\emptyset$. The set of tight filters $\widehat{E}_{\operatorname{tight}}$ is called the tight spectrum of $E$. It is the closure in $\hat{E}$ of the set of all ultrafilters (if $E$ is a Boolean algebra, then the tight filters are precisely the ultrafilters). For each $e\in E$, the set $Z_{e}:=\{\phi\in\widehat{E}_{\operatorname{tight}}:e\in\phi\}$ is compact open in $\widehat{E}_{\operatorname{tight}}$. For each $t\in S$, we have a homeomorphism $h_{t}:Z_{t^{*}t}\to Z_{tt^{*}}$ given by
@@
Let $u=\{u(s,t)\}_{s,t\in S}$ be a twist of the action $h:S\to\operatorname{PHomeo}(\widehat{E}_{{\operatorname{tight}}})$ described above. That is, $u(s,t)\in C_{u}(Z_{t^{*}s^{*}st})$, $t,s\in S$, are maps satisfying the conditions in [BE11, Definition 4.1], or in [BKM25, Definition 3.4]. Also let $\mathcal{L}_{u}$ be the twist associated to $u$, as in [BKM25, Subsection 4.3]. Then $(\mathcal{G}_{{\operatorname{tight}}}(S),\mathcal{L}_{u})$ is an ample twisted groupoid, and every ample twisted groupoid is of this form, see Remark 7.25 below.
We will now represent the twisted inverse semigroup $(S,u)$ via spatial partial isometries on the spaces $\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))$, $p\in[1,\infty]$. We recall that for any localizable measure $\mu$, in particular for a counting measure, we have a natural inverse semigroup $\operatorname{SPIso}(L^{p}(\mu))\subseteq B(L^{p}(\mu))$ of contractive weighted composition operators called spatial partial isometries on $L^{p}(\mu)$, see [BKM25, Definition 2.26], [Phi12, Definition 6.4], [Gar21, Definition 6.2].
@@
Let $S$ be an inverse semigroup with zero and let $u=\{u(s,t)\}_{s,t\in S}$ be a twist of the associated action $h:S\to\operatorname{PHomeo}(\widehat{E}_{{\operatorname{tight}}})$. For each $p\in[1,\infty]$, we have a map $v:S\to\operatorname{SPIso}\big(\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))\big)$ into spatial isometries on $\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))$ given by
@@
where $\xi\in\ell^{p}(\mathcal{G})$, $\phi\in Z_{s^{*}s}$, $t,s\in S$. Denoting by $v_{t}^{*}$ the unique generalized inverse of $v_{t}$ in $\operatorname{SPIso}\big(\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))\big)$, we get
@@
is a Banach algebra naturally isometrically isomorphic to the twisted reduced groupoid algebra $F_{\operatorname{r}}^{p}(\mathcal{G}_{{\operatorname{tight}}}(S),\mathcal{L}_{u})$ via an isomorphism that restricts to $\operatorname{\overline{span}}\{v_{t}v_{t}^{*}:t\in S\}\cong C_{0}(\widehat{E}_{\operatorname{tight}})$.
@@
By [BKM25, Remark 5.2], we have an isometric representation $\pi:C_{0}(\widehat{E}_{\operatorname{tight}})\to B\big(\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))\big)$ given by the formula $\pi(a)\xi[s,x]=a\big(h_{s}(x)\big)\xi([s,x])$ and then the pair $(\pi,v)$ forms a covariant representation of the twisted action that integrates to the regular representation of $F(\mathcal{G}_{{\operatorname{tight}}}(S),\mathcal{L}_{u})$ on $\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))$. In particular, $v_{t}$ is a spatial partial isometry, with the generalized inverse being $v_{t}^{*}=v_{t^{*}}\pi(\overline{u(t,t^{*})})$, and $v_{t}v_{t}^{*}=\pi(1_{Z_{tt^{*}}})$ for each $t\in S$. This implies that we have an isometric isomorphism $C_{0}(\widehat{E}_{\operatorname{tight}})\cong\pi(C_{0}(\widehat{E}_{\operatorname{tight}}))=\operatorname{\overline{span}}\{v_{t}v_{t}^{*}:t\in S\}$ that extends to an isometric isomorphism $F_{\operatorname{r}}^{p}(\mathcal{G}_{{\operatorname{tight}}}(S),\mathcal{L}_{u})\cong F^{p}_{\operatorname{r},{\operatorname{tight}}}(S,u)$ induced by the integrated form $\pi\rtimes v$ of the covariant representation $(\pi,v)$. ∎
@@
When the twist $u$ is trivial, we write $F^{p}_{\operatorname{r},{\operatorname{tight}}}(S)$ for $F^{p}_{\operatorname{r},{\operatorname{tight}}}(S,u)$. Then the map $v$ in Proposition 7.23 is a tight representation of $S$ in the sense of [BKM25, Definition 6.12]. In particular, $F^{p}_{\operatorname{r},{\operatorname{tight}}}(S)$ is a reduced version of the tight $L^{p}$-operator algebra $F^{p}_{\operatorname{tight}}(S)$ of $S$ introduced in [BKM25, Definition 6.14]: $v$ induces a natural representation $F^{p}_{\operatorname{tight}}(S)\to F^{p}_{\operatorname{r},{\operatorname{tight}}}(S)$ (which, by [GL17, Theorem 6.19], is an isometric isomorphism if $S$ is countable and the groupoid $\mathcal{G}_{{\operatorname{tight}}}(S)$ is amenable).
@@
By [BKM25, Corollary 6.18 and Lemma 4.23], every twisted ample groupoid $(\mathcal{G},\mathcal{L})$ is isomorphic to $(\mathcal{G}_{{\operatorname{tight}}}(S),\mathcal{L}_{u})$ for some inverse semigroup $S$ (one may take $S$ to be the collection of all compact open bisection in $S(\mathcal{L})$) and some twist $u$. Thus the algebras $F^{p}_{\operatorname{r},{\operatorname{tight}}}(S,u)$ introduced in Proposition 7.23 model all reduced twisted groupoid $L^{p}$-operator algebras associated to ample groupoids.
@@
Retain the notation and assumptions from Proposition 7.23 and fix $p\in[1,\infty]$. Denote by $\mathcal{G}_{\operatorname{H}}(S)$ the set of Hausdorff points in $\mathcal{G}_{{\operatorname{tight}}}(S)$. Then the obvious contractive projection $P_{\operatorname{H}}$ from $\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))$ onto $\ell^{p}(\mathcal{G}_{\operatorname{H}}(S))\subseteq\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))$ commutes with elements of $F^{p}_{\operatorname{r},{\operatorname{tight}}}(S,u)$, and if $S$ is countable, then the compressed Banach algebra
@@
Let $P\subseteq[1,\infty]$ be a non-empty set. We denote the Banach algebra generated by the $\ell^{\infty}$-direct sum of representations $S\to\operatorname{SPIso}\big(\ell^{p}(\mathcal{G}_{{\operatorname{tight}}}(S))\big)$, $p\in P$, from Proposition 7.23, by $F^{P}_{\operatorname{r},{\operatorname{tight}}}(S,u)$. We denote the compression of $F^{P}_{\operatorname{r},{\operatorname{tight}}}(S,u)$ to the subspace $\bigoplus_{p\in P}\ell^{p}(\mathcal{G}_{\operatorname{H}}(S))$ by $F^{P}_{\operatorname{e},{\operatorname{tight}}}(S,u)$. We also write $F^{P}_{{\operatorname{tight}}}(S,u):=F^{P}(\mathcal{G}_{{\operatorname{tight}}}(S),\mathcal{L}_{u})$ and skip writing $u$ when $u$ is trivial.
@@
In Definition 7.28 we slightly changed the terminology from [EP16]. Namely, the authors of [EP16] call weakly-fixed what we call fixed, and they call fixed what we call trivially fixed (our naming is consistent with [EP16, Definition 4.1], though). Local contractiveness is as in [EP16, Definition 6.1], closedness and minimality appear as conditions in [EP16, Theorems 3.16 and 5.5], respectively. Topological freeness is our invention and it improves a condition appearing in [EP16, Theorem 4.10], see Remark 7.32 below.
@@
Let $S$ be an inverse semigroup with zero. Then $S$ is closed (resp. topologically free or minimal) if and only if $\mathcal{G}_{\operatorname{tight}}(S)$ is Hausdorff (resp. topologically free or minimal). If $S$ is locally contractive, then $\mathcal{G}_{\operatorname{tight}}(S)$ is.
@@
The claims about $S$ being closed, minimal, and locally contractive follow by [EP16, Theorems 3.16, 5.5 and 6.5]. For the claim about topological freeness, recall that topological freeness of $\mathcal{G}_{\operatorname{tight}}(S)$ is equivalent to topological freeness of $h$ in the sense of [KM21, Definition 2.20], by [KM21, Lemma 2.23]. This means that for every $t\in S$, the set $\{\phi\in Z_{t^{*}t}:h_{t}(\phi)=\phi\}\setminus\bigcup_{f\in F_{t}}Z_{f}$ has empty interior. By [EP16, Proposition 2.5] and density of ultrafilters in $\widehat{E}_{\operatorname{tight}}$, every non-empty open set in $\widehat{E}_{\operatorname{tight}}$ contains a set of the form $Z_{e}$, for some $e\in E$. Thus topological freeness of $\mathcal{G}_{\operatorname{tight}}(S)$ is equivalent to the condition that for any $t\in S$ and $e\in E$, if $Z_{e}\subseteq\{\phi\in Z_{t^{*}t}:h_{t}(\phi)=\phi\}$, then $Z_{e}\cap\bigcup_{f\in F_{t}}Z_{f}\neq\emptyset$. In view of Lemma 7.30 and the fact that $Z_{e}\cap Z_{f}\neq\emptyset$ if and only if $ef\neq 0$, this in turn is equivalent to saying that for any $t\in S$ and $e\in E$, if $e$ is fixed by $t$, then $ef\neq 0$ for some $f\in F_{t}$, which is exactly what we called topological freeness of $S$. ∎
@@
The condition (ii) in [EP16, Theorem 4.10] tries to characterize effectiveness of $\mathcal{G}_{\operatorname{tight}}(S)$ in terms of $S$. It says that for every $e\in E$ fixed by $t\in S$, $e$ can be covered by elements in $F_{t}$. When $\mathcal{G}_{\operatorname{tight}}(S)$ is Hausdorff (equivalently $S$ is closed), this condition is equivalent to $S$ being topologically free, but in general topological freeness is is weaker. This agrees with the fact that effectiveness is equivalent to topological freeness for Hausdorff groupoids, but in general is stronger.